# Let the Void Be Void: Robust Open-Set Semi-Supervised Learning via Selective Non-Alignment

### v3

@@ Introduction
Semi-supervised learning (SSL) improves performance by leveraging unlabeled data, typically assuming that all unlabeled samples belong strictly to one of the known in-distribution (ID) classes (Sohn et al. 2020). However, practical unlabeled datasets inevitably form open sets, containing unknown out-of-distribution (OOD) samples. To address this limitation, Open-Set Semi-Supervised Learning (OSSL) learns from open-set unlabeled data to simultaneously improve ID classification and reliably detect OOD samples.
However, generalizable OOD detection via OSSL remains challenging. This is primarily because the true OOD space is inherently vast and heterogeneous, whereas training provides only limited, potentially misleading glimpses of OOD samples. Recent state-of-the-art methods take opposite stances toward OOD: aggressive or conservative. On the aggressive side, SCOMatch (Wang et al. 2024) collapses the diverse OOD universe into a single synthetic class, causing geometric collapse and overconfidence on seen OOD samples. Similarly, ProSub (Wallin et al. 2024b) creates a broad ID subspace spanning all ID clusters, aggressively pulling or pushing every unlabeled sample toward or away from this subspace even when pseudo-labels are noisy. Consequently, unseen OOD samples that land inside the broad ID subspace—yet outside any specific ID cluster—are misclassified as ID. On the conservative side, SSB (Fan et al. 2023) trains per-class one-vs-all (OVA) classifiers but discards low-confidence unlabeled data, losing valuable geometric cues that could improve feature space discrimination between ID and novel OOD.
@@
Early approaches to OSSL identify and exclude OOD samples, thus minimizing their negative impact on ID classifiers (Yu et al. 2020; Guo et al. 2020; Saito, Kim, and Saenko 2021). Recent methods, however, have shifted toward actively incorporating OOD samples. To leverage representations from OOD data, auxiliary self-supervised tasks, such as orientation prediction (Huang et al. 2021) or feature consistency enforcement (Wallin et al. 2024a), have been explored. However, these general-purpose tasks do not directly target OOD detection, resulting in limited gains.
@@
$\nabla_{z_{i}}\mathcal{L}_{\text{USNA}}=\frac{1}{T\|z_{i}\|}\big{(}I-\hat{z}\hat{z}^{\top}\big{)}\Big{(}\sum_{j=1}^{K}\alpha_{j}\,\hat{\mu}_{j}-\Phi_{i}\,\hat{\mu}_{\hat{k}}\Big{)}$ (3)
@@
This design selectively aligns only confident ID samples ($\Phi_{i}=1$) toward their respective class prototypes. In contrast, uncertain or OOD samples ($\Phi_{i}=0$) receive no targeted pull and instead undergo a non-selective angular update determined by the averaged term $\sum_{j}\alpha_{j}\hat{\mu}_{j}$, preventing unwanted alignment to any specific prototype. Complementing the $CC$ and $OD$ heads, SNA addresses a key limitation of their objectives (e.g., cross-entropy or pseudo-label consistency), which may unintentionally enlarge feature norms even for incorrectly pseudo-labeled samples. SNA’s stricter dual-gate mechanism classifies these ambiguous samples as uncertain ($\Phi_{i}=0$). The corresponding $\mathcal{L}_{\text{USNA}}$ then provides an orthogonal corrective “torque,” redirecting uncertain features away from incorrectly matched prototypes without aligning them elsewhere. Consequently, norm growth occurs exclusively along reliable directions, effectively preventing uncertain samples from accumulating large norms in misleading regions. Jointly, SNA, $CC$, and $OD$, yield a more discriminative and robust feature space, suitable for challenging open-set scenarios.
@@
To further assess discriminability, we analyze feature norms and prototype alignment. Figure 2(b) reveals a clear norm gap between ground-truth ID and OOD samples. This separation reflects how USNA encourages reliable samples to align with prototypes—thereby accumulating norm through $CC$ and ODOD—while preventing misaligned or unlabeled samples from doing so. As a result, ID samples exhibit significantly higher norms than both seen and unseen OODs. Figure 2(c) illustrates the average cosine similarity between each sample category and the ID prototypes. ID samples exhibit high similarity with their respective prototypes, whereas both seen and unseen OODs consistently display low similarity to all prototypes. Collectively, SNA, together with $CC$ and $OD$, naturally shapes a geometric structure consisting of dense, high-norm ID “galaxies” clearly separated by a sparse, low-norm OOD “interstellar void,” effectively generalizing to unseen OOD distributions.
@@
OOD detector. We follow the loss formulation of OpenMatch (Saito, Kim, and Saenko 2021) and SSB (Fan et al. 2023) to train the OOD detector. The total loss includes the One-vs-All (OVA) loss $\mathcal{L}_{\text{od}}^{\text{OVA}}$ for labeled data, an entropy minimization loss $\mathcal{L}_{\text{od}}^{\text{em}}$ to sharpen predictions on unlabeled samples, and a soft open-set consistency regularization loss $\mathcal{L}_{\text{od}}^{\text{SOCR}}$ that promotes prediction consistency under augmentation. In addition, we incorporate a pseudo-negative loss $\mathcal{L}_{\text{od}}^{\text{neg}}$, which leverages confidently identified OOD samples from $\mathcal{U}$ to enhance ID-OOD separation. To further improve robustness, both weak and strong augmentations are applied to these negative samples. The final OOD detector loss is defined as:
@@
Each $\lambda$ controls the relative importance of its corresponding loss component.
@@
Datasets and Backbones. For CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), we use WideResNet-28-2 (Zagoruyko 2016) and vary the number of ID/OOD classes with 25 or 50 labeled samples per class. For CIFAR-10, six animal classes are selected as ID classes, and the remaining four are used as OOD. In CIFAR-100, we first group classes by their semantic super-classes and then assign them into disjoint ID and OOD subsets. We evaluate two settings with different OOD proportions: one with 55 ID and 45 OOD classes, and another with 80 ID and 20 OOD classes. For each configuration, 25 or 50 labeled samples per ID class are used for training. For ImageNet30 (Hendrycks and Gimpel 2016), a 30-class subset of ImageNet (Deng et al. 2009), we adopt ResNet-18 (He et al. 2016) and use 20 classes as ID and 10 as OOD, with 5% of the ID data labeled (65 samples per class). For TinyImageNet (Le and Yang 2015), we use WideResNet-28-4 on a 100 ID / 100 OOD split, with 10% labeled data per ID class.
Baselines. We compare against recent OSSL methods, including IOMatch, SSB, ProSub, and SCOMatch. All methods are evaluated under the same dataset splits and hardware environment (H100 or RTX 3090), using official or faithfully reproduced implementations.
@@
We analyze the impact of varying the OOD confidence threshold $\eta_{\text{ID}}$ in our dual-gate ID selection, with the classifier threshold fixed at $\tau_{\text{ID}}=0.99$ and softmax temperature 0.5. All nonzero thresholds outperform the classifier-only gating ($\eta_{\text{ID}}=0$), demonstrating the importance of incorporating the OOD detector for reliable ID selection. On both CIFAR-10 and CIFAR-100, performance steadily improves as $\eta_{\text{ID}}$ increases from 0 and remains stable around moderate values (e.g., 0.5–0.7). While overly conservative settings (e.g., $\eta_{\text{ID}}=0.9$) may slightly reduce accuracy by filtering out some useful ID samples, this behavior reflects the binary nature of ID/OOD separation and indicates that our method performs robustly across a wide range of threshold values.
@@ A. Analysis of Feature Dynamics in SNA
To support our claim that Selective Non-Alignment (SNA) promotes robust feature structuring by deliberately delaying norm growth for uncertain samples, we conduct a gradient-based analysis of the feature dynamics. Unlike conventional contrastive methods that align all samples toward class prototypes, SNA explicitly prevents alignment for uncertain and OOD samples ($\Phi_{i}=0$), applying only angular repulsion without directional attraction. This selective non-alignment complements the classifier ($\mathcal{L}_{x}$, $\mathcal{L}_{u}$) and OVA detector ($\mathcal{L}_{\text{od}}^{\text{OVA}}$), mitigating their tendency to amplify norm even for noisy or misclassified features.
@@ A.4 SNA Loss: Angular Update
Here, $\hat{\mu}_{j}$ is the normalized prototype of class $j$, $\alpha_{j}$ is the soft pseudo-label probability for class $j$, and $\Phi_{i}\in\{0,1\}$ is the dual-gate confidence mask.
This gradient is orthogonal to $\hat{z}_{i}$, since:
@@
which implies it is purely angular and thus does not alter the norm $\|z_{i}\|$.
@@
Because the projection MLP $\mathcal{P}$ is smooth and approximately preserves direction in early training, the tangential property of $\nabla_{z_{i}}\mathcal{L}_{\text{USNA}}$ is partially retained after backpropagation to $f$. That is, $\nabla_{f}\mathcal{L}_{\text{USNA}}$ primarily induces a rotational (re-orienting) force rather than radial expansion, leading to directional adjustment without increasing $\|f\|$.
@@ A.5 Combined Dynamics and the Role of SNA
  • This purely angular update in the projected space $z_{i}=\mathcal{P}(f_{i})$ is backpropagated through the projection Jacobian $J_{\mathcal{P}}=\frac{\partial z}{\partial f}$, inducing a tangential force on the backbone feature $f$ that does not promote norm growth. This component acts as a corrective signal that can partially cancel the radial pull from the consistency loss $\mathcal{L}_{u}$.
@@
  • Furthermore, the repulsion term $\sum_{j}\alpha_{j}\hat{\mu}_{j}$ in the SNA loss applies a uniform angular push away from all class prototypes. This prevents $z_{i}$ from aligning with any specific prototype, maintaining a low $\cos\theta_{k}$ for all $k$. Since Closed-set Classifier and OOD Detector losses amplify norm in proportion to $\cos\theta_{k}$, this misalignment suppresses their norm-growing effect:
@@
As a result, norm growth for uncertain or OOD features is suppressed, and such embeddings remain in a low-norm, diffuse region. This mechanism is essential for maintaining a clear separation between ID and OOD samples, as illustrated in Figure 2.
@@
A.6 Conclusion
The SNA module introduces an angular corrective signal that selectively suppresses norm growth for uncertain samples. While classifier-based losses (e.g., cross-entropy, consistency, and OVA) typically promote norm growth via radial attraction—even for mislabeled or ambiguous samples—USNA applies only angular repulsion without contributing to radial expansion. The resulting feature geometry is well-structured: confident ID samples converge into compact, high-norm clusters, while uncertain and OOD samples remain dispersed in low-norm regions. This contrastive separation enhances robustness against pseudo-label noise and strengthens generalization to unseen OODs.

### v4

@@ Introduction
Semi-supervised learning (SSL) improves performance by leveraging unlabeled data, typically assuming that all unlabeled samples belong strictly to one of the known in-distribution (ID) classes (Sohn et al. 2020). However, practical unlabeled datasets inevitably form open sets, containing unknown out-of-distribution (OOD) samples. To address this limitation, Open-Set Semi-Supervised Learning (OSSL) learns from open-set unlabeled data to simultaneously improve ID classification and reliably detect OOD samples, including OOD classes present in the unlabeled pool (seen OOD) and novel OOD classes that appear only at test time (unseen OOD).
However, generalizable OOD detection via OSSL remains challenging. This is primarily because the true OOD space is inherently vast and heterogeneous, whereas training provides only limited, potentially misleading glimpses of OOD samples. Recent state-of-the-art methods take opposite stances toward OOD: aggressive or conservative. On the aggressive side, SCOMatch (Wang et al. 2024) collapses the diverse OOD universe into a single synthetic class, causing geometric collapse and overconfidence on seen OOD samples. Similarly, ProSub (Wallin et al. 2024b) creates a broad ID subspace spanning all ID clusters, aggressively pulling or pushing every unlabeled sample toward or away from this subspace even when pseudo-labels are noisy. Consequently, unseen OOD samples that fall within the broad ID subspace but outside any specific ID cluster tend to be misclassified as ID. On the conservative side, SSB (Fan et al. 2023) trains per-class one-vs-all (OVA) classifiers but discards low-confidence unlabeled data, losing valuable geometric cues that could improve feature space discrimination between ID and novel OOD.
@@
Early approaches to OSSL identify and exclude OOD samples, thus minimizing their negative impact on ID classifiers (Yu et al. 2020; Guo et al. 2020; Saito et al. 2021). Recent methods, however, have shifted toward actively incorporating OOD samples. To leverage representations from OOD data, auxiliary self-supervised tasks, such as orientation prediction (Huang et al. 2021) or feature consistency enforcement (Wallin et al. 2024a), have been explored. However, these general-purpose tasks do not directly target OOD detection, resulting in limited gains.
@@
[…]
@@
This design selectively aligns only confident ID samples ($\Phi_{i}=1$) toward their respective class prototypes. In contrast, uncertain or OOD samples ($\Phi_{i}=0$) receive no targeted pull and instead undergo a non-selective angular update determined by the averaged term $\sum_{j}\alpha_{j}\hat{\mu}_{j}$, preventing unwanted alignment to any specific prototype. Complementing the $CC$ and $OD$ heads, SNA addresses a key limitation of their objectives (e.g., cross-entropy or pseudo-label consistency), which may unintentionally enlarge feature norms even for incorrectly pseudo-labeled samples. SNA’s stricter dual-gate mechanism classifies these ambiguous samples as uncertain ($\Phi_{i}=0$). The corresponding $\mathcal{L}_{\text{USNA}}$ then provides an orthogonal corrective “torque”, redirecting uncertain features away from incorrectly matched prototypes without aligning them elsewhere. Consequently, norm growth occurs exclusively along reliable directions, effectively mitigating the risk of uncertain samples from accumulating large norms in misleading regions. Jointly, SNA, $CC$, and $OD$, yield a more discriminative and robust feature space, suitable for challenging open-set scenarios.
@@
To further assess discriminability, we analyze feature norms and prototype alignment. Figure 2(b) reveals a clear norm gap between ground-truth ID and OOD samples. This separation reflects how USNA encourages reliable samples to align with prototypes—thereby accumulating norm through the synergistic pull of $CC$, ODOD, and SNA alignment—while applying a selective repulsion to misaligned or unlabeled samples. As a result, ID samples exhibit significantly higher norms than both seen and unseen OODs. Figure 2(c) illustrates the average cosine similarity between each sample category and the ID prototypes. ID samples exhibit high similarity with their respective prototypes, whereas both seen and unseen OODs consistently display low similarity to all prototypes. Collectively, SNA, together with $CC$ and $OD$, naturally shapes a geometric structure consisting of dense, high-norm ID “galaxies” clearly separated by a sparse, low-norm OOD “interstellar void”, effectively generalizing to unseen OOD distributions.
@@
OOD detector. We follow the loss formulation of OpenMatch (Saito et al. 2021) and SSB (Fan et al. 2023) to train the OOD detector. The total loss includes the One-vs-All (OVA) loss $\mathcal{L}_{\text{od}}^{\text{OVA}}$ for labeled data, an entropy minimization loss $\mathcal{L}_{\text{od}}^{\text{em}}$ to sharpen predictions on unlabeled samples, and a soft open-set consistency regularization loss $\mathcal{L}_{\text{od}}^{\text{SOCR}}$ that promotes prediction consistency under augmentation. In addition, we incorporate a pseudo-negative loss $\mathcal{L}_{\text{od}}^{\text{neg}}$, which leverages confidently identified OOD samples from $\mathcal{U}$ to enhance ID-OOD separation. To further improve robustness, both weak and strong augmentations are applied to these negative samples. The final OOD detector loss is defined as:
@@
Each $\lambda$ weights the corresponding loss component.
@@
Datasets and Backbones. For CIFAR-10 and CIFAR-100 (Krizhevsky et al. 2009), we use WideResNet-28-2 (Zagoruyko 2016) and vary the number of ID/OOD classes with 25 or 50 labeled samples per class. For CIFAR-10, six animal classes are selected as ID classes, and the remaining four are used as OOD. In CIFAR-100, we first group classes by their semantic super-classes and then assign them into disjoint ID and OOD subsets. We evaluate two settings with different OOD proportions: one with 55 ID and 45 OOD classes, and another with 80 ID and 20 OOD classes. For each configuration, 25 or 50 labeled samples per ID class are used for training. For ImageNet30 (Hendrycks and Gimpel 2016), a 30-class subset of ImageNet (Deng et al. 2009), we adopt ResNet-18 (He et al. 2016) and use 20 classes as ID and 10 as OOD, with 5% of the ID data labeled (65 samples per class). For TinyImageNet (Le and Yang 2015), we use WideResNet-28-4 on a 100 ID / 100 OOD split, with 10% labeled data per ID class.
@@
Baselines. We compare SkipAlign against recent OSSL methods, including IOMatch, SSB, ProSub, and SCOMatch. All methods are evaluated under the same dataset splits and hardware environment (H100 or RTX 3090), using official or faithfully reproduced implementations.
@@
We analyze the impact of varying the OOD confidence threshold $\eta_{\text{ID}}$ in our dual-gate ID selection, with the classifier threshold fixed at $\tau_{\text{ID}}=0.99$ and softmax temperature 0.5 (Table 4). All nonzero thresholds outperform the classifier-only gating ($\eta_{\text{ID}}=0$), demonstrating the importance of incorporating the OOD detector for reliable ID selection. On both CIFAR-10 and CIFAR-100, performance steadily improves as $\eta_{\text{ID}}$ increases from 0 and remains stable around moderate values (e.g., 0.5–0.7). While overly conservative settings (e.g., $\eta_{\text{ID}}=0.9$) may slightly reduce accuracy by filtering out some useful ID samples, this behavior reflects the binary nature of ID/OOD separation and indicates that our method performs robustly across a wide range of threshold values.
@@ Conclusion
Acknowledgments
This work was supported by the Information and Communications Technology Planning and Evaluation (IITP) grant funded by the Korea Ministry of Science and ICT (MSIT) (Grant No. 2710085245); the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) through the Engineering Research Center (ERC) program (Grant No. 2710077975); and by the Mobile eXperience (MX) Business, Samsung Electronics Co., Ltd. Hyung-Sin Kim is the corresponding author.
@@ A. Analysis of Feature Dynamics in SNA
To support our claim that Selective Non-Alignment (SNA) promotes robust feature structuring by deliberately redirecting the feature dynamics for uncertain samples, we conduct a gradient-based analysis of the feature dynamics. Unlike conventional contrastive methods that align all samples toward class prototypes, SNA explicitly prevents alignment for uncertain and OOD samples ($\Phi_{i}=0$), applying only angular repulsion without directional attraction. This selective non-alignment complements the classifier ($\mathcal{L}_{x}$, $\mathcal{L}_{u}$) and OVA detector ($\mathcal{L}_{\text{od}}^{\text{OVA}}$), mitigating their tendency to amplify norm even for noisy or misclassified features.
@@ A.4 SNA Loss: Angular Update
Here, $\hat{\mu}_{j}$ is the normalized prototype of class $j$, $\alpha_{j}$ is the soft pseudo-label probability for class $j$, and $\Phi_{i}\in\{0,1\}$ is the dual-gate confidence mask. This gradient is orthogonal to $\hat{z}_{i}$, since:
@@
which implies it is purely angular.
@@
Because the projection MLP $\mathcal{P}$ is smooth and approximately preserves direction in early training, the tangential property of $\nabla_{z_{i}}\mathcal{L}_{\text{USNA}}$ is partially retained after backpropagation to $f$. That is, $\nabla_{f}\mathcal{L}_{\text{USNA}}$ primarily induces a rotational (re-orienting) force rather than radial expansion, leading to directional adjustment that steers the trajectory of $\|f\|$.
@@ A.5 Combined Dynamics and the Role of SNA
  • This purely angular update in the projected space $z_{i}=\mathcal{P}(f_{i})$ is backpropagated through the projection Jacobian $J_{\mathcal{P}}=\frac{\partial z}{\partial f}$, inducing a tangential force on the backbone feature $f$ that redirects this norm growth.
    This component acts as a corrective signal that can partially cancel the radial pull from the consistency loss $\mathcal{L}_{u}$.
@@
  • Furthermore, the repulsion term $\sum_{j}\alpha_{j}\hat{\mu}_{j}$ in the SNA loss applies a uniform angular push away from all class prototypes. This prevents $z_{i}$ from aligning with any specific prototype, maintaining a low $\cos\theta_{k}$ for all $k$. Since Closed-set Classifier and OOD Detector losses amplify norm in proportion to $\cos\theta_{k}$,
    this misalignment redirects their norm-growing effect away from prototypes:
@@
As a result, norm growth for uncertain or OOD features is applied in a repulsive direction , and such embeddings remain in a relatively low-norm, diffuse region. This mechanism is essential for maintaining a clear separation between ID and OOD samples, as illustrated in Figure 2.
@@
A.6 Theoretical Analysis of Generalization
In this section, we theoretically analyze the generalization properties of our proposed approach versus the conventional (K+1)-way classification method for OOD detection. To formalize this comparison, we utilize Rademacher complexity.
Theorem 1. Let $\mathcal{H}_{K+1}$ be a hypothesis space of functions $h:\mathcal{X}\rightarrow\mathbb{R}^{K+1}$ for $(K+1)$-way classification, where the $(K+1)$-th class is a synthetic representation for all Out-of-Distribution (OOD) data. Let $\mathcal{L}_{CE}$ be the standard multi-class cross-entropy loss. For any hypothesis $h\in\mathcal{H}_{K+1}$, with probability at least $1-\delta$ over a sample S of size m, the true risk $\mathcal{R}(h)$ is bounded by:
$\mathcal{R}(h)\leq\hat{\mathcal{R}}_{S}(h)+2\mathfrak{R}_{m}(\mathcal{L}_{CE}\circ\mathcal{H}_{K+1})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{m}})$
where $\hat{\mathcal{R}}_{S}(h)$ is the empirical risk and $\mathfrak{R}_{m}$ is the Rademacher complexity.
The Empirical Risk Minimization (ERM) principle for the $(K+1)$-way model compels the learning algorithm to map the inherently heterogeneous and expansive OOD data distribution, $P_{OOD}$, into a single, compact representation corresponding to the $(K+1)$-th class. This “aggressive alignment” results in a “geometric collapse” of the OOD space. Consequently, the hypothesis space $\mathcal{H}_{K+1}$ must possess a high capacity to represent the complex and non-linear decision boundaries required to separate the $K$ In-Distribution (ID) clusters from this artificially constructed OOD cluster. A hypothesis space of high capacity inherently exhibits a large Rademacher complexity, $\mathfrak{R}_{m}(\mathcal{L}_{CE}\circ\mathcal{H}_{K+1})$. This large complexity term yields a loose generalization bound, offering weak guarantees on the model’s performance, particularly for unseen OOD samples that do not conform to the collapsed geometry learned from seen OODs.
[…]
$\mathcal{R}(g\circ f_{\theta})\leq\hat{\mathcal{R}}_{S,\gamma}(g\circ f_{\theta})+\mathcal{O}(\frac{R_{feat}}{\gamma_{SNA}\sqrt{m}}+\sqrt{\frac{\log(1/\delta)}{m}})$
where $\hat{\mathcal{R}}_{S,\gamma}$ is the empirical margin loss.
The SNA framework reframes the OSSL problem from direct classification to representation learning. Its objective is to engineer a feature space that maximizes a composite geometric margin, $\gamma_{SNA}$, composed of both norm-based and angle-based separation. The gradient dynamics of the SNA loss, $\mathcal{L}_{SNA}$, mechanistically achieve this.
As discussed in Appendix A.5, confident ID samples receive synergistic alignment forces (from $\mathcal{L}_{CC}$, $\mathcal{L}_{OD}$, and $\mathcal{L}_{SNA}$), leading to rapid norm growth in the correct class direction. Conversely, uncertain or OOD samples receive a selective angular repulsion (from $\mathcal{L}_{USNA}$). As established in Appendix A.4 , this repulsion just induces slow norm growth, and critically, it does so in a direction away from all ID prototypes.
This dynamic—rapid, aligned norm growth for ID vs. slower, repulsive norm growth for OOD—is precisely what maximizes the geometric margin $\gamma_{SNA}$. According to margin-based learning theory, the existence of a large margin $\gamma_{SNA}$ permits the downstream OOD detection task to be solved by a low-capacity hypothesis space $\mathcal{G}_{\gamma}$ (e.g., simple linear classifiers). The Rademacher complexity of such a space is inversely proportional to the margin. Since the SNA algorithm is explicitly designed to maximize $\gamma_{SNA}$, it concurrently minimizes the complexity term of the generalization bound. This results in a significantly tighter bound, ensuring robust generalization to unseen OOD data.
A.7 Conclusion
The SNA module introduces an angular corrective signal that selectively redirects the feature dynamics and resulting norm growth for uncertain samples. While classifier-based losses (e.g., cross-entropy, consistency, and OVA) typically promote norm growth via radial attraction—even for mislabeled or ambiguous samples—USNA applies angular repulsion that primarily steers uncertain features away from ID clusters. The resulting feature geometry is well-structured: confident ID samples converge into compact, high-norm clusters, while uncertain and OOD samples remain dispersed in low-norm regions. This contrastive separation enhances robustness against pseudo-label noise and strengthens generalization to unseen OODs.